OpenStudy (anonymous):

Write a linear factorization of the function. f(x) = x^4 + 81x^2

OpenStudy (ranga):

\[\Large a^2 + b^2 = (a + bi)(a - bi)\]

OpenStudy (anonymous):

So it's f(x) = x2 (x + 9i)(x - 9i)?

OpenStudy (ranga):

You may have to write x^2 as x * x because x^2 is not a linear factor.

OpenStudy (anonymous):

Thank you!

OpenStudy (ranga):

you are welcome.